IN 1987, joy was 3 times as old as ali .if ali was born 8 year after joy what age will each be in 1999?
Answer by ikleyn(52794)

.
In 1987, Joy was 3 times as old as Ali and 8 ears older than Ali.


From it, you may conclude that in 1987 Ali was 4 years old, while Joy was 12 years old.


Hence, in 1999, Ali was  4 + (1999-1987) = 4 + 12 = 16 years old,    ANSWER

          while Joy was 12 + (1999-1987) = 12 + 12 = 24 years old.   ANSWER

Solved.

It is mental solution, without using equation.

If you want an algebraic solution, here it is.


Let A be the age of Ali in 1987.

Then Joy's age at that time was  3A, according to the condition.


Also, we know that


3A - A = 8.      ("Ali was born 8 year after Joy")


So,


2A = 8  ====>  A = 8/2 = 4 years (Ali's age in 1987).


From this point, you can complete the solution by the same way as I did it before.


You will get the same answer :  Ali was 16 years old in 1999;  Joy was 24 years old at that time.

Solved (two times with two methods).